Briana is a certain age. Carlee is 4 years older than Briana, and Dauna is 3 years older than Carlee. Together the sum of their ages is 44.
I have b for briana's age
b+4 for Carlee's age
b+7 for Dauna's age
Now I am stuck as to how to write the equation....
44=b+4+b+7+b ???????? please help

Answer by ankor@dixie-net.com(22740)

Briana is a certain age. Carlee is 4 years older than Briana, and Dauna is 3 years older than Carlee. Together the sum of their ages is 44.
I have b for briana's age
b+4 for Carlee's age
b+7 for Dauna's age
Now I am stuck as to how to write the equation....
44=b+4+b+7+b ???????? please help
:
Like you have:
b + 4 + b + 7 + b = 44
combine like terms
3b + 11 = 44
subtract 11 from both sides
3b = 44 - 11
3b = 33
divide both sides by 3
b = 33/3
b = 11 is Briana's age
then
11 + 4 = 15 is Carlees age
and
11 + 7 = 18 is Dauna's age
;
;
Check: 11 + 4 + 18 = 44